### v1

Abstract
Shadows significantly impact computer vision tasks, particularly in outdoor environments. State-of-the-art shadow removal methods are typically too computationally intensive for real-time image processing on edge hardware. We propose ShadowRemovalNet, a novel method designed for real-time image processing on resource-constrained hardware. ShadowRemovalNet achieves significantly higher frame rates compared to existing methods, making it suitable for real-time computer vision pipelines like those used in field robotics. Beyond speed, ShadowRemovalNet offers advantages in efficiency and simplicity, as it does not require a separate shadow mask during inference. ShadowRemovalNet also addresses challenges associated with Generative Adversarial Networks (GANs) for shadow removal, including artefacts, inaccurate mask estimations, and inconsistent supervision between shadow and boundary pixels. To address these limitations, we introduce a novel loss function that substantially reduces shadow removal errors. ShadowRemovalNet’s efficiency and straightforwardness make it a robust and effective solution for real-time shadow removal in outdoor robotics and edge computing applications.
keywords:
[…]
1 Introduction
Shadows, a common physical phenomenon observed in natural images and videos, are obscure regions where direct light from a source is either entirely or partly obstructed by an object. As depicted in Fig. 1, a shadow is typically divided into two parts: the umbra and the penumbra. The umbra represents the region of the shadow where all light is obstructed, while the penumbra is an area where only a portion of light is obstructed by the object. While shadows can provide valuable visual information, such as the shape and depth of objects, they can also pose challenges in various computer vision and image analysis applications. For instance, in the object segmentation task, the edges of shadows may be mistaken for actual object edges. Similarly, in tracking scenarios, shadows can lead to tracking of incorrect targets. Therefore, the removal of shadows from images and videos is a crucial step for improving the performance of these tasks. By eliminating shadows, we can reduce potential errors and improve the accuracy of computer vision applications.
Figure 1: An illustration of shadow formation as it can occur in field robotics. Part a) showcases a real-world scenario where a tractor, equipped with a weed spraying system, is traversing across a field. The tractor, obstructing the sunlight, casts a shadow on the ground, obscuring the presence of weeds and plants. Part b) offers a schematic diagram that elucidates the scientific concept of shadow formation. It comprises three main components: an object (represented by the tractor), an illumination source (symbolised by the Sun), and a surface. The interplay between these elements results in the casting of a shadow on the surface. The shadow manifests distinct regions, each labelled for better understanding: "Non-shadow region", "Penumbra", "Umbra", and "Shadow region". The "non-shadow region" is the area that receives direct light from the illumination source. The shadow itself is bifurcated into two zones: the umbra and the penumbra. The umbra, being the core of the shadow, is its darkest part where the object completely obstructs the light. Conversely, the penumbra is a peripheral lighter region where only a portion of the light is blocked by the object. This figure serves as an effective tool for visualising and comprehending the intricate process of shadow formation, setting the stage for our discussion on shadow removal techniques for robotised weed spot spraying
The context of shadow removal becomes particularly crucial when applied to field robotics. Computer vision methods are commonly used in field robotics for object detection, classification, localisation, and more. However, changing lighting conditions due to cloud cover, the movement of the robot relative to the sun, and shadows created by objects in the environment all complicate the computer vision task. For instance, shadows cast by an agricultural vehicle (Fig. 1a) create areas of reduced visibility. Therefore, the removal of shadows from images and videos is helpful as an image preprocessing or normalisation step. By effectively eliminating shadows, we could enhance the accuracy of computer vision methods used in field robotics, thereby increasing their performance.
The study of shadow removal has seen significant advancements in recent years, primarily due to the application of deep learning (DL)-based methods. DL-based methods have leveraged large-scale datasets to predict shadow masks for shadow removal, formulate linear illumination models, and apply networks to predict correspondent shadow parameters. Notable contributions in this field include the direction-aware method by (Hu et al., 2020), which uses spatial image context for shadow removal, and RIS-GAN by (Zhang et al., 2020)., which implements shadow removal in a coarse-to-fine manner. Other significant works include the context aggregation network by (Cun et al., 2020), and the boundary-aware network by (Niu et al., 2023). Recent innovations have further propelled this field. For instance, ShadowFormer by (Guo et al., 2023), the early shadow removal method based on transformer, and novel frameworks designed by ST-CGAN and ARGAN to jointly perform shadow detection and removal.
Supervised methods are fundamentally limited by the difficulty in obtaining large-scale paired datasets. A paired dataset is one where the same scene is presented both with and without shadows. Therefore, unsupervised DL-based methods have also emerged, yielding favourable results in performing shadow removal on unpaired datasets. An unpaired dataset is one without access to shadow-free images of each scene that appears in the training set. These methods must learn to remove shadows from unpaired datasets, which are generally easier to collect. Unsupervised learning algorithms work by discovering patterns and structures within the input data. In the context of shadow removal, these algorithms can learn to distinguish between shadowed and non-shadowed regions in an image, and then create an instance of the image without the shadows. One of the key challenges in unsupervised shadow removal is dealing with the inherent ambiguity of the task. Since shadows can vary greatly in terms of size, shape, colour, and intensity, it can be difficult for an algorithm to accurately identify and remove all shadows from an image. However, recent advancements in deep learning have led to the development of more sophisticated models that are capable of handling this complexity. For instance, some unsupervised methods leverage generative models such as Generative Adversarial Networks (GANs) (Zhang et al., 2020) to perform shadow removal. These models consist of two parts: a generator that creates images without shadows, and a discriminator that tries to distinguish between real images and the ones generated by the generator. Through this adversarial process, the generator learns to create images that are indistinguishable from real images without shadows.
Figure 2: A comparison of our proposed shadow removal method (ShadowRemovalNet) against state-of-the-art models, based on their PSNR performance (dB), inference time (S), frame per second (fps), number of parameters (M), and computational complexity in giga FLOPS (G). The models were tested on the ISTD dataset (Wang et al., 2018) using a 2080Ti GPU device, corresponding to the No. in Tab. 5. Each data point represents a different model, with its position indicating the model’s inference time (x-axis) and PSNR performance (y-axis). Additional dimensions of comparison, such as the number of parameters and computational complexity, provide a more comprehensive evaluation of each model’s efficiency and effectiveness. This visualization allows for an effective comparison of our method against other leading models in the field of shadow removal, taking into account not only performance but also model complexity and computational demands.
In the field of shadow removal, we have identified and addressed several critical challenges. The first challenge relates to the artifacts produced by Generative Adversarial Networks (GANs). An essential element of GANs, the shadow attention decoder, creates attention maps that significantly influence the performance of shadow removal. Moreover, one major challenge in network training is known to be the instability of adversarial training. Simultaneously, these algorithms are often trained in a supervised way on pairs of shadowed and shadow-free images of the same instances. To create these pairs, an initial image of a scene with shadows is taken, followed by another after the shadow-causing object has been removed. Furthermore, a binary mask that marks the existence of shadows must be annotated either manually or automatically if the model needs mask information. This procedure significantly restricts the number and diversity of scenes that can be collected and can lead to noisy or biased supervision. Shadow removal models trained with such data do not perform well with a wide range of real-world shadow images (Kang et al., 2023). Additionally, it is important to note that there is no guarantee that areas free of shadows will remain consistent under varying sunlight conditions. This uncertainty can compromise the reliability of the paired training images. The second challenge involves inaccurate mask estimates for shadow images in existing datasets. These estimates often contain errors, such as incorrect labels around the boundary or minor shadow objects. Even meticulously hand-crafted masks occasionally introduce coarse boundaries, which can impede the effectiveness of shadow removal methods.
To address the aforementioned challenges, we have adopted a two-pronged approach. First, we have introduced a novel loss function that specifically addresses the problem of unbalanced supervision within shadow boundary pixels and pixels inside shadows. This function enhances the supervision of shadow boundary pixels, thus significantly decreasing the errors caused by shadow removal. Second, to overcome the limitations posed by the scarcity of high-quality image pairs, we have collected a substantial dataset of 10,000 natural shadow-free images from the internet. This large-scale collection significantly enhances the robustness and generalization ability of our shadow removal method. This approach has led to superior performance in shadow removal tasks and demonstrates the potential of using internet-sourced, shadow-free images to improve the training of shadow removal algorithms.
As shown in Fig. 2, our method distinguishes itself through its efficiency, operating at a speed of 66 frames per second when running on an nVidia 2080Ti, and its simplicity, as it does not necessitate a mask during inference. This end-to-end approach demonstrates its practicality for real-time image processing. Existing state-of-the-art methods operate at only approximately 9 frames per second on the same hardware, which falls short of the required speed for real-time image processing. Acknowledging this limitation, we have engineered our method to operate at a significantly higher speed, thereby effectively addressing this issue. Beyond the advantage of speed, our method also surpasses ShadowFormer in terms of efficiency and simplicity. In contrast to ShadowFormer, our method simplifies the process by eliminating the need for a mask during inference and operates at a speed that is eight times faster. This combination of factors positions our method as an optimal choice for shadow removal.
We summarize our contributions as follows:
  • Addressing the challenge of artifacts produced by Generative Adversarial Networks (GANs).
  • Addressing the issue of inaccurate mask estimates for shadow images in existing datasets.
  • Introducing a novel loss function to counteract the problem of inconsistent supervision between pixels inside shadows and shadow boundary pixels.
  • Developing an efficient and straightforward method that operates at high speeds on commodity hardware and does not require a mask during inference.
These targeted improvements make our method a robust and efficient solution for shadow removal. The remainder of this paper is structured as follows: Section 2 provides a detailed presentation of our proposed approach. Section 3 outlines the experimental setup and implementation details of our method, evaluates our approach’s performance, and compares it with several state-of-the-art methods. Section 4 discusses the Strengths and limitations of our approach and future work. Finally, Section 5 summarises the main findings of this paper and highlights the potential impact of our work.
Figure 3: Schematic Diagram of the Proposed ShadowRemovalNet. This diagram illustrates the architecture of our proposed Shadow Removal Network (ShadowRemovalNet), designed to generate a variety of shadow-free images from a single input. The network consists of three key components: an encoder, a latent space, and a decoder. The encoder transforms the input image into a lower-dimensional representation. This output is then mapped onto a high-dimensional latent space. The decoder, in turn, generates a diverse set of shadow-free images from this latent space (red dots). The resulting shadow-free images are displayed as samples on the right side of the diagram.
2 Method
In this section, we present our method for shadow removal. As shown in Fig. 3, the ShadowRemovalNet model’s ability to generate multiple enhanced versions of an image is a key feature that sets it apart. This is achieved by using the Probabilistic Enhancement Module (PEM), which transforms the statistics of the received features using random samples from the prior and posterior distributions of the latent space. These distributions capture the diversity in the input image, allowing the model to generate multiple interpretations of the original image. Each interpretation, or enhanced version, represents a different possible understanding of the original image. This could include variations in lighting, colour balance, contrast, and other visual elements. This capability is particularly useful in complex environments where the same image can be interpreted in different ways. For example, in the context of agricultural field robotics, it allows the model to identify and highlight different potential areas of interest within the same image, such as different types of weeds or variations in plant health. By generating multiple enhanced versions of the image, the model provides a more comprehensive understanding of the scene. This approach also reduces the risk of missing important details that might be visible in one interpretation but not in others. A comprehensive explanation of our method is provided in the following sections.
2.1 Preliminary
Image shadow removal is the task of estimating a shadow-free image $Y=\{y_{1},\ldots,y_{N}\}$ from a shadow image $x=\{x_{1},\ldots,x_{N}\}$, where $y_{i}$ and $x_{i}$ refer to the respective pixel intensities. An observation from Fig. 1 indicates that the shadowed region receives less light compared to the non-shadowed region. This is due to the obstruction of direct light in the shadowed area, resulting in a decrease in light intensity. In the context of this task, the input to a conventional convolution layer is represented by feature maps ($X_{in}\in\mathbb{R}^{C\times H\times W}$), where $C$ is the number of input channels, and $H$ and $W$ are the height and width of the input feature maps respectively. The output ($X_{out}\in\mathbb{R}^{C^{\prime}\times H^{\prime}\times W^{\prime}}$) of the convolution operation is computed by convolving $X_{in}$ with a convolution weight ($W\in\mathbb{R}^{C^{\prime}\times C\times k\times k}$), where $C^{\prime}$ is the number of output channels and $k\times k$ is the size of the convolution kernel. Additionally, in some methods, a binary ground-truth shadow mask ($M\in\{0,1\}^{H\times W}$) is used. Each entry in the mask ($M_{ij}$) is either 1 or 0, indicating whether a pixel in the input $X_{in}$ is a shadow pixel or a non-shadow pixel respectively. In this study, we use unpaired data for training our model. This data consists of shadow images that do not have corresponding shadow-free images or binary ground-truth shadow masks. Consequently, we generate shadow images and the corresponding binary shadow masks, the process for which is described in the subsequent section.
2.2 Shadow Synthesis
The process of shadow synthesis begins with a shadow-free image, denoted $\bm{x}^{sf}$. The objective is to generate an image, $\bm{x}^{shade}\in\mathbb{R}^{H\times W\times 3}$, where all pixels are shaded and share similar attenuation properties. This is accomplished by applying the affine model proposed in (Inoue and Yamasaki, 2021) to compute $\bm{x}^{shade}$ from $\bm{x}^{sf}$.
The model is summarized by the following two equations that are mathematically equivalent:
$x^{sf}_{ijk}=\alpha_{k}+\gamma~{}x^{shade}_{ijk}$ (1)
$\iff$ $x^{shade}_{ijk}=\frac{1}{\gamma}~{}x^{sf}_{ijk}-\frac{\alpha_{k}}{\gamma},$ (2)
where $i$ and $j$ represent indices for the row and column axes, respectively, for all images. $\alpha_{k}$ and $\gamma$ as parameters representing a constant offset and a scaling factor, respectively.
Eq. 1 describes a relationship between a variable representing a non-shadow image ($x_{ijk}^{sf}$) and a variable representing a shaded image ($x_{ijk}^{shade}$), with $\alpha_{k}$ and $\gamma$ as parameters representing a constant offset and a scaling factor, respectively. Eq. 2 is the inverse of Eq. 1, suggesting that the shaded image can be derived from the non-shadow image by subtracting the offset and dividing by the scaling factor.
The final image, $\bm{x}^{s}$, which contains shadows in certain regions, is obtained by combining $\bm{x}^{sf}$ and $\bm{x}^{shade}$ using alpha composition. The shadow matte $\bm{m}$ serves as the alpha factor in this composition, represented by the following equation:
$x^{s}_{ijk}$ $=(1-m_{ij})x^{sf}_{ijk}+m_{ij}x^{shade}_{ijk},$ (3)
where $x_{ijk}^{s}$ denotes the pixel value at position $(i,j)$ and channel $k$ in the synthesized shadow image, $m_{ij}$ represents the matte value at position $(i,j)$ indicating the transparency of the shadow, and $x_{ijk}^{sf}$ and $x_{ijk}^{shade}$ represent the pixel values at the same position in the shadow-free and shaded images, respectively.
The pixel value in the synthesized shadow image is a weighted combination of the pixel values from the shadow-free and shaded images, with the weight determined by the matte value. If the matte value is 0 (indicating no shadow), the pixel value in the synthesized shadow image is equal to the pixel value in the shadow-free image. Conversely, if the matte value is 1 (indicating full shadow), the pixel value in the synthesized shadow image is equal to the pixel value in the shaded image.
The shadow matte $\bm{m}$ is a crucial component in the shadow synthesis process. It is derived from masks in the ISTD dataset (Wang et al., 2018) and serves as the alpha factor in the alpha composition of $\bm{x}^{sf}$ and $\bm{x}^{shade}$ to generate the final image $\bm{x}^{s}$ with shadows. In the shadow matte $\bm{m}$, the value of each pixel $m_{ij}$ represents the degree of shadow at that location. Specifically, $m_{ij}=1$ signifies that the pixel is within the umbra, the fully shaded inner region of a shadow where light from the light source is completely blocked by the occluding body. On the other hand, $0\leq m_{ij}\leq 1$ indicates that the pixel is in the penumbra, the partially shaded outer region of the shadow where only a portion of the light source is obscured by the occluding body. Lastly, $m_{ij}=0$ denotes that the pixel is not in the shadow, i.e., it is fully illuminated. This method allows for a realistic representation of shadows, taking into account the varying degrees of darkness in the umbra and penumbra regions, as well as the transition between shadowed and illuminated areas. The use of the shadow matte $\bm{m}$ in the alpha composition ensures that the shadows in the synthesized image $\bm{x}^{s}$ are consistent with the shadow-free image $\bm{x}^{sf}$ and the darkened image $\bm{x}^{shade}$, thereby enhancing the overall realism of the synthesized shadows.
Figure 4: demonstrates an example of mask dissociation, where the ground truth mask is separated into complementary components that capture coarse body and fine boundary details, respectively. (b) shows the body mask of the ground truth, with larger pixel values near the centre of the target. (c) shows the detail mask, with larger values near the object boundaries. The sum of the body mask (b) and detail mask (c) equals the complete ground truth mask (a).
2.3 Shadow Boundaries
In our research, we observed that the difficulty in predicting a pixel’s characteristics is closely tied to its position. Pixels near the edge, often affected by a cluttered background, are more likely to be mispredicted, while central pixels exhibit higher prediction accuracy due to the salient target’s internal consistency. To address this, we introduced a novel approach of mask dissociation, where we separate the original mask into two parts: a body mask and a detail mask, as shown in Fig. 4. This dissociation is achieved using Distance Transformation (DT), a conventional algorithm for image processing. The DT converts a binary image into a new image where each foreground pixel value represents its shortest Euclidean distance to the nearest background pixel. Thus, DT maps object silhouettes to topographic relief maps reflecting the spatial proximity to exterior boundaries. After the distance transformation, the pixel value in the new image, which we refer to as the body image, relies not just on whether it is part of the foreground or the background, but also on its relative position. Therefore, the body image represents the central pixels that are comparatively straightforward to predict.
We get the detail image by subtracting the body image from the original image.This detail image, which we use as the detail mask in subsequent experiments, primarily focuses on pixels distant from the body area. To eliminate background interference, we multiply the newly created masks with the ground-truth binary image. The distance transformation can be expressed as
$I^{{}^{\prime}}(p)$ =$\min\limits_{q\in I_{bg}}f(p,q),\qquad p\in I_{fg}\\ 0,\qquad p\in I_{bg}\end{aligned}\right.$ (4)
where, $I^{{}^{\prime}}(p)$ is the transformed image after applying the Distance Transformation (DT). $f(p,q)$ is the metric function used to measure the distance between pixels. $p$ is a pixel in the foreground of the image ($I_{fg}$). $q$ is the nearest pixel in the background of the image ($I_{bg}$) to pixel $p$.
The masks are generated as
$Mask\Rightarrow\left\{\begin{aligned} BM&=I*I^{{}^{\prime}}\\ DM&=I*(1-I^{{}^{\prime}})\end{aligned}\right.$ (5)
where $BM$ means the body mask and $DM$ represents the detail mask. Now the original mask has been dissociated into the body mask and detail mask. This dissociation allows our network to learn both the body and detail features with different characteristics respectively. This mask dissociation approach provides a more nuanced method for pixel prediction, taking into account the relative position of pixels and their respective prediction difficulties. This method enhances the learning process, enabling the network to effectively learn different features with varying characteristics as explained in the next section.
2.4 Network Structure
ShadowRemovalNet employs a two-branch U-Net-based feature extractor to map input images to representations, as illustrated in Fig. 3. These representations are subsequently fed into the Probabilistic Enhancement Module (PEM), which transforms the enhancement statistics of the input to create the enhanced image. The training branch of the feature extractor constructs posterior distributions using the raw original shadow image and its corresponding shadow-free image as inputs. Conversely, the test branch uses only a shadow image as input to estimate the prior distribution of a single raw shadow image. The PEM encodes the diversity in the input image, enabling ShadowRemovalNet to generate multiple enhanced versions of the image that capture the different possible interpretations of the original image. To achieve this, ShadowRemovalNet uses a prior/posterior block to build the distribution of possible enhancements. This block constructs both a mean and a standard deviation distribution, using ${1\times 1}$ convolutions to transform the input data matrix into a series of distributions that capture the diversity in the input image.
Probabilistic Enhancement Module (PEM). The proposed PEM is a key component of the ShadowRemovalNet model. The primary objective of ShadowRemovalNet is to eliminate shadows from images without compromising the quality of the original image. This is particularly beneficial in outdoor settings where lighting conditions can vary greatly, leading to the formation of shadows. Therefore, we use PEM to capture these properties.
The PEM method is an extension of the method proposed by (Nuriel et al., 2021) for the Adaptive Instance Normalization (AdaIN) algorithm. The AdaIN algorithm is a popular technique in image processing, particularly for style transfer tasks that adjust the feature statistics of a content image to match those of a style image. Specifically, it modifies the mean and standard deviation of content features to align with the style image. This shifts the visual attributes of the content image closer to the style reference, effectively blending aspects of both while preserving scene composition. The output inherits stylistic qualities from the style image while depicting the same content as the original. However, the PEM method introduces an element of randomness into this process. This allows the PEM method to perform style transformation for any image without relearning, making it particularly useful for applications like shadow removal in outdoor environments where the style of the image (i.e., lighting conditions) can vary greatly. The proposed method uses random samples drawn from the posterior distributions of the mean and standard deviation. These sampled statistics parameterize the adaptive instance normalization (AdaIN) operation for style transfer. By propagating stochasticity from the posterior into AdaIN, the framework models uncertainty to produce diverse stylizations that avoid deterministic solutions. The unique stylizations better explore the complex style manifold. This can be formulated as:
${\rm{PEM}}\left({\bf{x}}\right)={\bm{b}}\left({\frac{{{\bf{x}}-\bm{\mu}\left({\bf{x}}\right)}}{{\bm{\sigma}\left({\bf{x}}\right)}}}\right)+{\bm{a}},$ (6)
where $\bm{b}$ and $\bm{a}$ are two random samples from the posterior distributions of the mean and standard deviation, respectively.
This allows PEM to generalize the AdaIN algorithm and apply it to shadow removal without the need for known content and style images. The posterior distributions are described in the next section.
Training Stage. During the training stage, the shadow image and its corresponding shadow-free image are used to learn the posterior distributions of the latent space as follows:
${\bm{a}}\sim\mathcal{N}_{\rm{m}}\left({\bm{\mu}\left({{\bm{y}},{\bf{x}}}\right),\bm{\sigma}^{2}\left({{\bm{y}},{\bf{x}}}\right)}\right),$ (7)
where ${\bm{a}}$ is a random sample from the mean posterior distribution. $\mathcal{N}_{\rm{m}}$ represents the $N$-dimensional Gaussian distribution of the mean. ${\bm{\mu}\left({{\bm{y}},{\bf{x}}}\right)}$ is the mean of the distribution, which is a function of the input image (x) and the reference map (${\bm{y}}$). ${\bm{\sigma}^{2}\left({{\bm{y}},{\bf{x}}}\right)}$ is the variance of the distribution, which is also a function of the input image and the reference map.
${\bm{b}}\sim\mathcal{N}_{\rm{s}}\left({\bm{m}\left({{\bm{y}},{\bf{x}}}\right),\bm{v}^{2}\left({{\bm{y}},{\bf{x}}}\right)}\right),$ (8)
where ${\bm{b}}$ is a random sample from the standard deviation posterior distribution. $\mathcal{N}_{\rm{s}}$ represents the $N$-dimensional Gaussian distribution of the standard deviation. ${\bm{m}\left({{\bm{y}},{\bf{x}}}\right)}$ is the mean of the distribution. ${\bm{v}^{2}\left({{\bm{y}},{\bf{x}}}\right)}$ is the variance of the distribution.
Equations Eq. 7 and Eq. 8 are used in the training stage of the ShadowRemovalNet model to learn the posterior distributions of the latent space. The input image and its corresponding reference image are used to determine the mean and variance of the distribution. The random sample (${\bm{a}}$) and (${\bm{b}}$) are then used in the PEM to transform the statistics of the received features. This allows the model to generate multiple enhanced versions of the image that capture the different possible interpretations of the original image.
Testing Stage. In the testing stage, the latent space generated for PEM are determined only by the input image to learn the prior distributions of the latent space as follows:
${\bm{a}}\sim\mathcal{N}_{\rm{m}}\left({\bm{\mu}\left({{\bf{x}}}\right),\bm{\sigma}^{2}\left({{\bf{x}}}\right)}\right),$ (9)
where ${\bm{a}}$ is a random sample from the mean prior distribution. $\mathcal{N}_{\rm{m}}$ represents the ($N$)-dimensional Gaussian distribution of the mean. ${\bm{\mu}\left({{\bf{x}}}\right)}$ is the mean of the distribution. ${\bm{\sigma}^{2}\left({{\bf{x}}}\right)}$ is the variance of the distribution.
${\bm{b}}\sim\mathcal{N}_{\rm{s}}\left({\bm{m}\left({{\bf{x}}}\right),\bm{v}^{2}\left({{\bf{x}}}\right)}\right),$ (10)
where ${\bm{b}}$ is a random sample from the standard deviation prior distribution. $\mathcal{N}_{\rm{s}}$ represents the ($N$)-dimensional Gaussian distribution of the standard deviation. ${\bm{m}\left({{\bf{x}}}\right)}$ is the mean of the distribution. ${\bm{v}^{2}\left({{\bf{x}}}\right)}$ is the variance of the distribution.
Equations Eq. 9 and Eq. 10 are used in the testing stage of the ShadowRemovalNet model to learn the prior distributions of the latent space. The input image is used to determine the mean and variance of the distribution. The random sample (${\bm{a}}$) and (${\bm{b}}$) are then used in the PEM to transform the statistics of the received features. The ShadowRemovalNet model is applied multiple times to the same input image to generate multiple enhancement variants. This is done by re-evaluating only the PEM and the output block, without retraining the entire model, making ShadowRemovalNet very efficient. The resulting diverse enhancement samples are then used for Maximum Probability estimation that takes the enhancement sample with the maximum probability as the final estimation using Maximum A Posteriori (MAP). The MAP estimator is a Bayesian estimator that maximizes the posterior distribution to find the most probable value of the parameter given the data. It is often used when we have some prior knowledge about the parameter distribution.
2.5 Loss function
The training procedure for ShadowRemovalNet is a multi-step process with the primary objective of instructing the model to generate high-quality shadow-free images from input shadow images. This process also incorporates diversity into the shadow removal process. The initial phase of this process involves minimizing the variational lower bound, a standard practice in the training of Conditional Variational Autoencoder (cVAE) models. The subsequent phase is to find a significant embedding of illumination statistics in the latent space. This is achieved by deploying a posterior network. We introduce a network trained to infer posterior feature distributions after encoding a content-style pair. Specifically, it identifies latent content features and maps them to predicted posterior distributions over the mean and standard deviation parameters. By randomly sampling from these predicted distributions, we obtain unique stylization parameters for each forward pass. Our framework thus enhances the standard style transfer output by embedding variability based on the learned feature uncertainties. This produces a rich range of enhanced results formalized through the stochastic sampling process. The third phase addresses the problem of inconsistent supervision between pixels inside shadows and shadow boundary pixels. This imbalance can lead to substantial errors during the shadow removal process.
In the training process of ShadowRemovalNet, the PEM module plays a crucial role in predicting the shadow-free image. It does so by receiving random samples ${\bm{a}}$ and ${\bm{b}}$ from equations Eq. 7 and Eq. 8, respectively. The enhancement loss, denoted as $L_{\rm{e}}$, is then computed. This loss quantifies the discrepancies between the reference image and the predicted image, serving as a penalty to the model if the output drifted from the reference. The enhancement loss is calculated using the formula:
${L_{\rm{e}}}={L_{\rm{mse}}}+\lambda{L_{\rm{vgg16}}},$ (11)
where $L_{\rm{mse}}$ represents the mean square error loss and $L_{\rm{vgg16}}$ signifies the perceptual loss. The variable $\lambda$ is a weight parameter.
The mean square error loss $L_{\rm{mse}}$ and the perceptual loss $L_{\rm{vgg16}}$ are two prevalent metrics employed to assess the performance of shadow removal algorithms. The mean square error loss calculates the average squared difference between the predicted and reference images, while the perceptual loss, introduced by (Johnson et al., 2016), measures the differences between the high-level features of the predicted and reference images.
Beyond the minimization of the enhancement loss $L_{\rm{e}}$, the training process for ShadowRemovalNet also incorporates the use of Kullback-Leibler (KL) divergences $D_{\rm{KL}}$. These divergences serve to align the posterior distributions with the prior distributions, as shown in equations Eq. 12 and Eq. 13:
$L_{m}={D_{{\rm{KL}}}}\left({{\mathcal{N}_{\rm{m}}}\left({{\bf{x}}}\right)\left\|{{\mathcal{N}_{\rm{m}}}\left({{\bm{y}},{\bf{x}}}\right)}\right.}\right),$ (12)
$L_{s}={D_{{\rm{KL}}}}\left({{\mathcal{N}_{\rm{s}}}\left({{\bf{x}}}\right)\left\|{{\mathcal{N}_{\rm{s}}}\left({{\bm{y}},{\bf{x}}}\right)}\right.}\right),$ (13)
where $\bm{m}$ and ${\bm{s}}$ are the mean and the standard deviation, respectively.
The quality of shadow restoration, particularly at the shadow boundaries, is crucial in shadow removal. To enhance the supervision of pixels near the shadow boundaries in the final predicted shadow-free image, a boundary loss function is designed. As mentioned in Sec. 2.3, we introduced a novel approach of mask dissociation, where we separate the original mask into two parts: a body mask (BM) and a detail mask (DM), as shown in Fig. 4. Therefore, the boundary loss term is defined as follows:
$L_{b}=||\hat{x^{sf}}\odot\tilde{M}_{DM}-x^{sf}\odot\tilde{M}_{DM}||_{1},$ (14)
where $\hat{x^{sf}}$ is the predicted shadow-free image, $x^{sf}$ is the reference image, and $\tilde{M}_{DM}$ is the weighted detail shadow mask. The symbol ||.$||.||_{1}$ denotes the L1 norm, and $\odot$ represents the Hadamard product. This boundary loss term helps to ensure the quality of shadow removal, particularly at the shadow boundaries.
The total loss function employed for training ShadowRemovalNet is a weighted sum of the enhancement loss $L_{\rm{e}}$, the KL divergences $D_{\rm{KL}}$ between the posterior and prior distributions, and shadow boundaries loss:
$L=\alpha L_{e}+\beta(L_{m}+L_{s})+\gamma L_{b},$ (15)
In this Eq. 15, $\alpha$, $\beta$, and $\gamma$ are weight parameters that control the relative importance of the enhancement loss $L_{e}$, the KL divergences $(L_{m}+L_{s})$, and the boundary loss $L_{b}$ in the total loss function $L$. The specific values of these parameters used in the training process will depend on the characteristics of the dataset and the desired performance of the model. By minimizing this total loss function, the model can learn an effective mapping from the input degraded images to the corresponding enhanced images, while also aligning the posterior and prior distributions in the latent space and enhancing the supervision of pixels near the shadow boundaries. This allows the model to generate high-quality enhanced images while also incorporating diversity into the shadow removal process.
3 Experiments
In this section, we present the experiments conducted to evaluate the performance of our proposed method.
3.1 Experiment setup
Benchmark datasets. We used three benchmark datasets to carry out a series of shadow removal experiments: ISTD Dataset (Wang et al., 2018), the adjusted ISTD (ISTD+) Dataset (Le and Samaras, 2019), and SRD Dataset (Qu et al., 2017). We use the ISTD dataset, which is composed of 1330 training and 540 testing triplets. Each triplet includes shadow images, masks, and shadow-free images. This dataset provides a robust set of images that serve as the basis for our shadow removal algorithm training and testing. We also use the ISTD+ dataset, an enhanced version of the ISTD dataset. It incorporates an image processing algorithm to mitigate the illumination inconsistency between the shadow and shadow-free images found in the original ISTD dataset. The quantity of triplets in the ISTD+ dataset mirrors that in the ISTD dataset. Lastly, we use the SRD dataset, which includes 2680 training and 408 testing pairs of shadow and shadow-free images. It is important to note that this dataset does not provide ground truth shadow masks. To overcome this limitation, we incorporate the predicted masks furnished by DHAN (Cun, Pun, and Shi 2020) for our testing phase only.
Training datasets. The scarcity of high-quality image pairs, each consisting of a shadow and its corresponding shadow-free version of the same scene, poses a significant challenge to the development of robust shadow removal algorithms. The limited dataset size in the state-of-the-art SRD and ISTD datasets may not capture the diversity and complexity of real-world scenes, leading to suboptimal generalization and performance. To overcome these limitations, we collected a substantial dataset of 10,000 natural shadow-free images from the internet. This large-scale collection significantly enhances the robustness and generalization ability of our shadow removal method.
In our study, we trained two distinct models. The ’Model Standard’ was trained on the standard SRD and ISTD datasets, while the ’Model Plus’ was trained on the 10,000 natural shadow-free images collected from the internet. The performance and comparative analysis of these two models are reported in the subsequent section. This dual-model approach allowed us to evaluate the effectiveness of our method in different scenarios and further validate the benefits of our large-scale, internet-sourced dataset.
Evaluation metrics. We adopt three evaluation metrics to assess the performance of our shadow removal method. These metrics are applied in line with previous works (Qu et al., 2017; Wang et al., 2018; Hu et al., 2019; Guo et al., 2023; Zhu et al., 2022a). 1) Root Mean Square Error (RMSE): We use RMSE as a quantitative evaluation metric for the shadow removal results. The RMSE is computed in the CIE LAB colour space, which models perceptual uniformity across the luminance and chrominance channels. Lower RMSE values indicate better shadow removal, with the metric assessing errors over corresponding pixels in the prediction output and shadow-free target. 2) Peak Signal-to-Noise Ratio (PSNR): We also employ PSNR to evaluate the performance of different approaches in the RGB colour space. Higher PSNR values represent better results. 3) Structural Similarity (SSIM) (Wang et al., 2004): In addition to RMSE and PSNR, we adopt SSIM to further evaluate the shadow removal performance in the RGB color space. Similar to PSNR, higher SSIM values indicate better performance. Our evaluation is conducted on the original image resolutions and with a resolution of $256256x$ 256, following the approach of previous works (Jin et al., 2021; Zhu et al., 2022a; Fu et al., 2021). We compare our technique with various state-of-the-art methods on the ISTD (Wang et al., 2018), ISTD+ (Le and Samaras, 2019), and SRD (Qu et al., 2017) datasets in both quantitative and qualitative ways.
Implementation details. We’ve developed our shadow removal method using PyTorch 1.8, and it’s been tested on a Linux platform with NVIDIA RTX 2080Ti GPUs. Our training process involves random cropping of images into $256256x$ 256 patches. We’ve set the total training epochs to 500 and mini-batch size to 8 across all datasets. The Adam optimizer (Kingma and Ba, 2014) is used with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=1.0\times 10^{-08}$ and an initial learning rate of $lr_{initial}=1e-4$, which linearly decays to $lr_{final}=1e-6$ in the final $200$ epochs.
Figure 5: presents a side-by-side comparison of our method with state-of-the-art shadow removal methods on the ISTD dataset (Wang et al., 2018). The images are arranged from left to right as follows: the input image, methods by DHAN (Cun et al., 2020), Fu et al. (2021), Zhu et al. (2022b), ShadowFormer (Guo et al., 2023), our method, and the ground truth. Zoom in to see the details.
Figure 6: presents a side-by-side comparison of our method with state-of-the-art shadow removal methods on the SRD dataset (Qu et al., 2017). The images are arranged from left to right as follows: the input image, methods by DHAN (Cun et al., 2020), Fu et al. (2021), Zhu et al. (2022b), ShadowFormer (Guo et al., 2023), our method, and the ground truth. Zoom in to see the details.
Figure 7: Visual examples of error maps resulting from our method. The first column displays the original images with shadows. The second column shows the images with the shadows removed. The third column presents the predicted images with the shadows removed. The fourth column illustrates the error maps for the predictions, with white dots representing the error. The color scale on the right side of the error maps ranges from 0 to 255, indicating the magnitude of the error. Zoom in to see the details.
3.2 Comparison with State-of-the-Art Methods
In this section, we provide a comparable evaluation of our proposed method with various state-of-the-art (SOTA) shadow removal algorithms. These include one traditional method, namely Guo et al. (2012), and a number of deep learning-based methods such as MaskShadow-GAN (Hu et al., 2019), ST-CGAN (Wang et al., 2018), ARGAN (Ding et al., 2019), DSC (Hu et al., 2020), G2R (Liu et al., 2021), Fu et al. (2021), Jin et al. (2021), DHAN (Cun et al., 2020), Zhu et al. (2022b), SG-ShadowNet (Wan et al., 2022), BMNet (Zhu et al., 2022a), and ShadowFormer (Guo et al., 2023). Given the diversity in experimental settings adopted by prior shadow removal methods, we seek a fair comparison by conducting experiments on two primary settings over the ISTD dataset. These settings follow the methodologies of two recent works, Zhu et al. (2022b) and DHAN (Cun et al., 2020). The first setting involves resizing the results into a resolution of $256\times 256$ for evaluation. The second setting maintains the original image resolutions throughout both the training and testing stages. It is important to note that all shadow removal results from the competing methods are either quoted from the original papers, taken from the previous papers or reproduced using their official implementations. This ensures the integrity and fairness of our comparative analysis.
Table 1: The quantitative results of shadow removal using our models and recent method on ISTD (Wang et al., 2018) datasets. We put “-” to denote those models or results that are not available.
[…]
Quantitative Measure. The quantitative results on the testing sets over ISTD, ISTD+, and SRD are presented in Tables (Tab. 1, Tab. 2, and Tab. 3), respectively. Our proposed approach achieves superior quantitative performance across all datasets compared to existing methods, as evidenced by lower RMSE and higher PSNR and SSIM in shadow regions, non-shadow regions, and entire images. These consistent improvements demonstrate the strength of our shadow removal on diverse scenes. Interestingly, some existing methods degrade non-shadow regions post processing according to the non-shadow area (NS) results on the ISTD+ dataset in Tab. 3. By contrast, our approach enhances both shadow and non-shadow areas simultaneously. The ability to improve all image regions underscores the precision of our data-driven shadow manipulation techniques. Our method offers several advantages. It can effectively infer consensus point estimates to produce trace-less results by sampling multiple latent variables from the learned distribution. The proposed mask dissociation loss is designed to address the imbalance in supervision between shadow boundary pixels and pixels inside shadows. It does this by leveraging the contextual correlation between shadow and non-shadow regions. This approach allows for a more accurate and balanced supervision, leading to improved shadow detection and removal. In comparison to the most recent state-of-the-art methods, such as those proposed by (Ding et al., 2019) and Fu et al. (2021), which employed very deep and large backbones like U-Net256 and ResNeXt, our method is more efficient, as shown in Tab. 1. Despite using less than 3% of the number of parameters of these competitive approaches, our method delivers outstanding shadow removal performance on test sets.
Qualitative Measure.
Qualitative results on the ISTD (Fig. 5) and SRD (Fig. 6) datasets further showcase the advantages of our approach compared to other methods. The ISTD examples contain relatively simple scenes with high contextual similarity across images. Despite this, prior arts often generate illumination mismatches and artifacts along shadow boundaries in these samples. By contrast, our method produces results absent of tonal inconsistencies and edge effects. This holds even for subtle shadows in similar contexts that cause issues for existing techniques. We also demonstrate strong generalizability on the more complex and diverse SRD shadows. The ability to attenuate difficult shadows while preserving lighting coherence emphasizes the robustness of our data-driven framework. For instance, Fu et al. (2021) and DHAN (Cun et al., 2020) methods erroneously brighten some areas with poor lightness, such as the fourth row of Fig. 5, resulting in ghosting artifacts. Furthermore, almost all competing methods except ShadowFormer (Guo et al., 2023) fail to maintain the illumination coherencies among shadow and non-shadow areas, which severely disrupts the image structure and patterns, thereby generating sharp boundaries as shown in the fourth and fifth rows of Fig. 5. However, with the benefits of the proposed mask dissociation loss, It is evident that our techniques can effectively balance the supervision between shadow boundary pixels and pixels inside shadows. This improves shadow removal, even in complex lighting conditions and intricate shadow patterns. Furthermore, our approach enhances the overall image quality by preserving the integrity of non-shadow regions, thereby providing a more natural and visually pleasing result. On the contrary, the image structures on the SRD dataset are more complex and often feature various colours. In these SRD dataset samples, previous work has sometimes struggled to accurately remove shadows, particularly in areas with intricate structures or varied colours, such as the blue poster in the third row of Fig. 6. The complexity of the image structures and the diversity of colours can pose significant challenges for shadow removal algorithms, leading to less-than-optimal results. Despite the complexity and colour diversity, our method effectively removes shadows while preserving the original colours and structures. This results in high-quality images that maintain the integrity of the original scene, proving the effectiveness of our approach in handling complex and diverse image data.
Table 2: The quantitative results of shadow removal using our models and recent methods on SRD (Qu et al., 2017) datasets. We put “-” to denote those models or results that are not available.
[…]
Table 3: Quantitative comparison of our method with the state-of-the-art methods on ISTD+ datasets (Le and Samaras, 2019).
[…]
3.3 Generalization Ability
To validate the generalization ability of our proposed approach, we conducted experiments on the SBU-Timelapse dataset (Le and Samaras, 2022). This dataset is unpaired, meaning it does not include shadow-free images as ground truth. Therefore, we adopted non-reference evaluation indicators: Gradient Structure Similarity (NRSS) and Natural Image Quality Evaluator (NIQE) (Li et al., 2022). The NRSS was used to assess the naturalism of the created shadow-free images, while the NIQE was used to measure the perceptual similarity. For the NRSS metric, a higher value indicates better performance, meaning the denoised image is more similar to the original noisy image. For the NIQE metric, a lower value is indeed better, indicating that the image has a higher perceptual quality. We compared our method with several state-of-the-art methods that do not require masks during inference, including SID (Le and Samaras, 2022), DHAN (Cun et al., 2020), and SG-ShadowNet (Wan et al., 2022). As shown in Table 4, our approach surpasses the other three approaches across all metrics. When compared to SG-ShadowNet (Wan et al., 2022), our method managed to reduce the NIQE from 7.6532 to 3.7532, and achieved an increase in NRSS from 0.9324 to 0.9632. Our method also exhibited better performance when compared to SID (Le and Samaras, 2022) and DHAN (Cun et al., 2020). As depicted in Figure 8, our method achieves reasonable perceptual performance in complicated environments, thereby demonstrating the generalization ability of our method. These results highlight the effectiveness of our proposed method in handling complex and diverse image data.
Table 4: Quantitative comparison of our method with the state-of-the-art methods on the SBU-Timelapse dataset (Le and Samaras, 2022). $\uparrow$” indicates the higher the better and $\downarrow$” indicates the lower the better.
[…]
Figure 8: presents a side-by-side comparison of our method with state-of-the-art shadow removal methods on the SBU-Timelapse dataset (Le and Samaras, 2022). The images are arranged from left to right as follows: the input image, methods by SID (Le and Samaras, 2022), DHAN (Cun et al., 2020), SG-ShadowNet (Wan et al., 2022), our method. It is important to note that the SBU-Timelapse dataset is an unpaired dataset, meaning it does not include shadow-free images as ground truth. This aspect adds an additional layer of complexity to the task of shadow removal and serves as a testament to the effectiveness of the compared methods. Zoom in to see the details.
3.4 Computational efficiency
In this section, we analyse the computational efficiency of our proposed shadow removal method in comparison with several state-of-the-art models. The comparison is based on key metrics including frames per second (FPS), inference time, computational complexity in giga FLOPs (FLOPs), and the number of parameters (Params). All models were tested on the ISTD dataset using a 2080Ti GPU device. As shown in Table 5 and Fig. 2, our method significantly outperforms the other models in terms of FPS, achieving a rate of 66. This is considerably higher than the next best model, SP+M-Net, which achieves an FPS of 10. Our method also boasts the shortest inference time of 0.015 s, demonstrating its efficiency. In terms of computational complexity, our method requires only 7.5 giga FLOPs, which is significantly lower than the other models. For instance, the DHAN model requires 262.9 giga FLOPs, which is over 35 times more than our method. Furthermore, our method uses only 16.1M parameters, which is less than most of the other models. The ShadowFormer model uses slightly fewer parameters (9.3M), but it achieves a lower FPS of 8 compared to our method’s 66. In conclusion, our method not only delivers superior performance but also demonstrates impressive computational efficiency, making it a highly effective solution for shadow removal tasks for field robotics.
Table 5: Performance and Complexity Analysis of Shadow Removal Methods: A comparison of our method with state-of-the-art models on key metrics including PSNR performance, inference time, frames per second (FPS), number of parameters (Params), and computational complexity (FLOPs). All models were tested on the ISTD dataset using a 2080Ti GPU device.
Method FPS Time FLOPs PSNR Params DHAN (Cun et al., 2020) 9 0.117 262.9 29.12 21.8 SP+M-Net (Le and Samaras, 2019) 10 0.105 160.1 25.08 141.2 G2R (Liu et al., 2021) 4 0.254 113.9 25.07 113.9 Fu et al. (2021) 7 0.155 160.3 27.19 143.0 ShadowFormer (Guo et al., 2023) 8 0.127 100.9 32.21 9.3 Ours 66 0.015 7.5 32.76 16.1
Figure 9: Visual Enhancement in Robotic Weed Spot-Spraying: A Comparative Study of Images Before and After Shadow Removal
3.5 Application of Shadow Removal in Field Robotics
Our shadow removal method presents substantial benefits for two critical areas of field robotics: real-time image processing and image processing during dataset labelling.
3.5.1 Real-time Image Processing
Field robots often operate in dynamic environments with varying lighting conditions. This can significantly impact the uniformity of captured images, hindering the performance of computer vision pipelines within the robot. Our shadow removal method addresses this challenge by enabling real-time image processing. By removing shadows from the captured frames, the method ensures consistent image quality and improves the clarity of crucial features for object detection and recognition tasks. This is particularly valuable in agricultural settings where environmental lighting fluctuates throughout the day. Consistent image quality allows the robot’s computer vision system to perform more accurate object detection and recognition tasks, leading to improved decision-making and overall performance.
3.5.2 Image Processing during Dataset Labelling
Training robust and accurate deep learning models is highly dependent on the quality of the labelled data sets used. Human experts typically label these datasets by manually identifying and annotating objects of interest within the images. Our method can be applied during this labelling process to preprocess the images before they are presented to human labellers. By removing shadows, the method enhances the visibility of key features within the image. This reduces the potential for human error during labeling and ultimately leads to a higher quality labeled dataset. A well-labeled dataset is essential for training deep learning models that can generalize well to unseen data and perform effectively in real-world scenarios.
Figure 9 presents real-world images captured by a weed control robot, before and after shadow removal. The clear improvement in image quality and visibility of the weeds demonstrates the effectiveness of our method in improving the labelling process.
Tables 6 and 7 present the measured inference times (ms) and frame rates (fps) of the shadow algorithm, respectively. These tables quantify the performance of the algorithm on a Jetson-Orin AGX platform in various configurations. They explore the impact of input size (256x256 and 512x512 pixels) and weight precision (float32, float16, int8) for both PyTorch and TensorRT implementations. Analyzing these results is crucial for selecting the optimal configuration for real-world applications, striking a balance between speed and accuracy requirements.
Environment Weight Precision Input Size
512$\times$512 px 256$\times$256 px
PyTorch float32 80.87 $\pm$ 2.08 22.07 $\pm$ 0.42
TensorRT float32 78.32 $\pm$ 0.25 21.52 $\pm$ 0.36
TensorRT float16 51.79 $\pm$ 0.26 20.25 $\pm$ 0.26
TensorRT int8 46.96 $\pm$ 0.34 18.80 $\pm$ 0.24
Table 6: Inference times (ms) of the shadow algorithm measured on a Jetson Orin AGX embedded GPU platform varying input size (512$\times$512 and 256$\times$256 px) and weight precision (float32, float16, int8) for PyTorch and TensorRT implementations.
Environment Weight Precision Input Size
512$\times$512 px 256$\times$256 px
PyTorch float32 12.37 $\pm$ 0.31 45.33 $\pm$ 0.84
TensorRT float32 12.77 $\pm$ 0.04 46.47 $\pm$ 0.77
TensorRT float16 19.31 $\pm$ 0.10 49.39 $\pm$ 0.63
TensorRT int8 21.30 $\pm$ 0.15 53.21 $\pm$ 0.65
Table 7: Frame rates (fps) of the shadow algorithm measured on a Jetson Orin AGX embedded GPU platform varying input size (512$\times$512 and 256$\times$256 px) and weight precision (float32, float16, int8) for PyTorch and TensorRT implementations.
3.6 Ablation Study
An ablation study was conducted to understand the contribution of each component of the loss function to the performance of the ShadowRemovalNet model. The loss function is composed of three main components: the enhancement loss $L_{e}$, the KL divergence $(L_{m}+L_{s})$, and the boundary loss $L_{b}$. Each of these components is weighted by a corresponding weight parameter ($\alpha$, $\beta$, and $\gamma$ respectively) in the total loss function:
$L=\alpha L_{e}+\beta(L_{m}+L_{s})+\gamma L_{b}$
In the ablation study, each component of the loss function was removed one at a time, and the performance of the model was evaluated without that component. This allowed us to quantify the impact of each component on the overall performance of the model. The results of the ablation study, Tab. 8, showed that each component of the loss function contributes significantly to the performance of the ShadowRemovalNet model. The enhancement loss $L_{e}$ was found to be crucial for generating high-quality shadow-free images, while the KL divergence $(L_{m}+L_{s})$ played a key role in aligning the posterior and prior distributions in the latent space. The boundary loss $L_{b}$ was particularly important for enhancing the supervision of pixels near the shadow boundaries, leading to more accurate shadow removal in these areas.
Table 8: Contributions of the components of the loss function to the performance of the ShadowRemovalNet model tested on ISTD (Wang et al., 2018) datasets. .
[…]
4 Discussion
In this section, we discuss the implications of our findings, the strengths and weaknesses of our method, and potential avenues for future research.
Implications. The findings from our study have significant implications for field robotics. Our method for shadow removal using unpaired data demonstrates that it is possible to effectively identify and remove shadows from images, even without a direct comparison to shadow-free ground truth images. This opens up new possibilities for improving the accuracy and efficiency of computer vision in various outdoor lighting conditions. The use of unpaired data also means that our method can be applied in a wider range of scenarios, as it does not require the collection of paired shadow and shadow-free images. This could potentially simplify the data collection process and make our method more accessible for use in different environments and lighting conditions. Moreover, our method’s ability to handle complex and diverse image data suggests that it could be applied in a wide variety of field robotics applications, where challenges with lighting and shadows may exist.
Relationship with Unsupervised Method. Shadow image degradation is complex and irreversible, making it infeasible to directly supervise a shadow removal model with ground truth data. Therefore, unsupervised methods are better suited for this task. Rather than developing a new unsupervised approach, our work focuses on effectively utilizing synthetic labels from existing shadow removal datasets. While not perfect, these synthetic reference maps of shadow-free images can be easily obtained. We propose a framework that leverages conditional variational autoencoders and our proposed module to learn an enhancement model from these imperfect references. In summary, both our approach and unsupervised methods aim to train a shadow removal model without ground truth data. However, our framework provides a novel perspective, i.e., transforming the problem into distribution estimation and consensus building using readily available, but biased references. This represents a distinct approach from existing unsupervised techniques.
Training Dataset and Model Development. The development of robust shadow removal algorithms is significantly influenced by the quality and quantity of available training data. Currently, the scarcity of high-quality image pairs, each consisting of a shadow and its corresponding shadow-free version of the same scene, poses a considerable challenge. The largest available datasets for shadow removal, such as the state-of-the-art SRD and ISTD datasets, contain only around 100 unique scenes. This limited dataset size may not capture the diversity and complexity of real-world scenes, leading to suboptimal generalization to unseen scenes and inferior performance in shadow removal tasks. Models trained on such limited datasets may struggle to understand the variations in lighting conditions, scene composition, and other factors that affect the appearance of shadows. This could result in a failure to maintain consistent colors between the shadow and shadow-free regions, leading to unnatural or inconsistent results in shadow removal. The difficulty in collecting a large dataset with paired shadow/shadow-free images, given the various types of shadow masks and scenes, further compounds these challenges. To overcome these limitations, we have collected a substantial dataset of 10,000 natural shadow-free images from the internet. This large-scale collection significantly enhances the robustness and generalization ability of our shadow removal method. By synthesizing shadow images from this extensive dataset, we can train our shadow removal algorithms more effectively, thereby achieving superior performance in shadow removal tasks. This approach demonstrates the potential of using internet-sourced, shadow-free images to improve the training of shadow removal algorithms.
Strengths. Our method for shadow removal has several key strengths. Firstly, it effectively handles the inherent complexity of shadow removal, delivering high-quality results even in challenging lighting conditions. This is particularly important in outdoor settings, where lighting conditions can vary greatly. Secondly, our method does not require paired shadow and shadow-free images, which simplifies the data collection process and makes the method more versatile and easier to apply in different scenarios. This is a significant advantage, as collecting paired images can be time-consuming and impractical in many situations. Thirdly, our method demonstrates impressive computational efficiency. It outperforms several state-of-the-art models in terms of key performance metrics, making it a highly effective solution for real-time shadow removal tasks. Finally, our method shows strong generalization ability. It performs well on diverse image data, suggesting that it could be adapted for use in other related fields or applications. This opens up exciting possibilities for future research and development.
Limitations. Despite its strengths, our method also has some limitations. One of the main limitations is that it may struggle with extremely complex shadows or lighting conditions. While our method performs well under a variety of conditions, there may be certain scenarios where the complexity of the shadows exceeds the capabilities of our method. Another limitation is the reliance on unpaired data. While this allows for greater flexibility in data collection, it also means that our method does not have a direct comparison to a shadow-free ground truth image. This could potentially impact the accuracy of the shadow removal process. Finally, while our method has demonstrated strong performance on the datasets we tested, its performance on other datasets or real-world scenarios remains to be seen. It is possible that our method may not perform as well on different types of images or in different environments.
Future Work. There are several potential avenues for future research based on our current work. One of the primary areas of focus could be to improve the handling of extremely complex shadows or lighting conditions. This could involve refining our current method or exploring new techniques and approaches for shadow detection and removal. Another area for future work could be to test our method on a wider range of datasets and real-world scenarios. This would provide valuable insights into the generalizability and practical applicability of our method, and could also help to identify any potential limitations or areas for improvement that are not apparent from the current datasets. Additionally, future research could explore ways to optimize the computational efficiency of our method. This could involve refining the current algorithm, exploring more efficient algorithms, or leveraging advances in hardware technology.
5 Conclusions
Our study presents a novel method for shadow removal using unpaired data. Our method effectively handles the inherent complexity of shadow removal, delivering high-quality results even in challenging lighting conditions. Despite the complexity of the task, our method demonstrates impressive computational efficiency, outperforming several state-of-the-art models in terms of key performance metrics. However, like any method, ours is not without its limitations. It may struggle with extremely complex shadows or lighting conditions, and its performance on other datasets or real-world scenarios remains to be seen. Despite these limitations, our method opens up new possibilities for improving the accuracy and efficiency of outdoor computer vision in various lighting conditions. Looking ahead, there are several potential avenues for future research. These include improving the handling of complex shadows, testing our method on a wider range of datasets, optimizing computational efficiency, and exploring applications in other related fields. Through continued research and development, we believe our method has the potential to make significant contributions to computer vision for field robotics.
5.1 CO2 Emission Related to Experiments
[…]
Data Availability
Data will be made available on request
Acknowledgement
This research is funded by the partnership between the Australian Government’s Reef Trust and the Great Barrier Reef Foundation.
Additional Information
Competing interests The authors declare no competing interests.
[…]